# Li Tana,b , Shengrong Wanga,c

### v1

@@ Abstract
In this article, we propose a tamed theta Euler-Maruyama (EM) scheme for superlinearly growing neutral McKean-Vlasov stochastic differential delay equations driven by fractional Brownian motion with Hurst exponent $H\in(1/2,1)$. The analytical properties including uniqueness and existence, boundedness of moment and propagation of chaos are investigated. Moreover, the convergence rate of the numerical scheme is established.
Keywords :Neutral McKean-Vlasov SDEs; superlinearly growing; fractional Brownian motion; strong convergence rate; tamed theta EM
@@
Stochastic differential equations (SDEs) with coefficients that depend on the law of the solution were first studied by McKean [1], who was inspired by Kac’s programme in Kinetic Theory [2]. Subsequently, they have become a research hotspot. Bossy and Talay [3, 4] chose to approximate the McKean-Vlasov limit by replicating the behavior with a system of $n$ weakly interacting particles. They also studied the numerical approximation of the solutions to the McKean-Vlasov SDEs. The existence and uniqueness of McKean-Vlasov SDEs were discussed in [5, 6]. Lacker [7] investigated the well-posedness and the propagation of chaos for McKean-Vlasov SDEs. Since the explicit solutions of McKean-Vlasov SDEs are difficult to obtain, many types of McKean-Vlasov SDEs can be approximated by implicit Euler-Maruyama (EM) scheme [8, 16], adaptive EM scheme [9, 11], tamed EM scheme [10, 14], truncated EM scheme [15] and so on. They approximated the solutions of McKean-Vlasov SDEs using numerical methods whose coefficients are globally (or locally) Lipschitz continuous or exhibit super-linear growth.
Recently, there has been a tendency for research to focus on the McKean-Vlasov SDEs driven by fractional Brownian motion. The form is
@@
where the coefficients $b:\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d},\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ are Borel-measurable, here, $\mathcal{P}_{q}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite $q$-th moment. $\mathcal{L}(X_{t})$ is the law of $X$ at $t$ and $B^{H}=\{B_{t}^{H},t\in[0,T]\}$ is a fractional Brownian motion with Hurst exponent $H\in(0,1)$ defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, $B^{H}$ is a centered Gaussian process with the covariance function
$R_{H}(t,s)=\mathbb{E}(B_{t}^{H}B_{s}^{H})=\frac{1}{2}(t^{2H}+s^{2H}-|t-s|^{2H}),\ \ \forall t,s\in[0,T].$
When $H=1/2$, the fractional Brownian motion $\{B_{t}^{H}\}_{t\geq 0}$ corresponds to a standard Bromnian motion. The self-similarity and long memory properties of fractional Brownian motion make it applicable in various fields, such as atmospheric sciences, financial engineering and others. From then on, scholars conducted studies on (1.1) separately. He et al. [17] established the theory of propagation of chaos and studied the strong convergence of the EM method when the coefficients of (1.1) satisfy the globally Lipschitz condition.
@@
However, few studies considered the case where the coefficients of (1.1) grow super-linearly. Shen [18] demonstrated that, under certain averaging conditions, the solutions of McKean-Vlasov SDEs driven by fractional Brownian motion can be approximated by the solutions of the associated averaged McKean-Vlasov SDEs driven by fractional Brownian motion, in the sense of the mean square convergence, where the coefficients exhibit super-linear growth. Wang et al. [19] and Gao et al. [20] focus on the EM scheme for delayed McKean-Vlasov SDEs driven by fractional Brownian motion, whose drift coefficient contains a polynomial delay term.
Actually, in order to describe the impact of the past, stochastic delay differential equations (SDDEs) and neutral stochastic differential delay equations (NSDDEs) play a significant role in stochastic calculus, describing a wide variety of natural and man-made systems (see, e.g., [21, 22, 23]). The classical EM method is no longer applicable, as the coefficients of SDEs do not exhibit linear growth. It is known that the tamed EM shceme can be used to deal with superlinearly growing coefficients. Deng et al. [24] proposed two types of explicit tamed EM schemes for NSDDEs with superlinearly growing drift and diffusion coefficients and proved strong convergence for each of them, respectively. Cui et al. [25] studied the tamed EM methods to approximate the solutions for a type of of McKean-Vlasov NSDDEs in which the drift coefficients exhibit superlinear growth, and obtained the rate of convergence. We can refer to [23, 27] for more detail about tamed EM scheme in NSDDEs. It is worth noting that Tan and Yuan [28] modified the tamed theta EM scheme and showed the convergence of the numerical solutions, which provides us with research ideas for conducting this paper.
Drawing inspiration from the tamed approach introduced in [28] and previous models of McKean-Vlasov SDEs driven by fractional Brownian motion, we consider a tamed theta EM scheme for MV-NSDDEs driven by fractional Brownian motion with coefficients of superlinear growth, as an extension of (1.1). The form is
@@
where $\tau>0$ is the delay.
The structure of the rest of the paper is as follows. In Section 2, the mathematical preliminaries on the McKean-Vlasov NSDDEs with superlinear growth driven by fractional Brownian motion are presented. In Section 3, existence and uniqueness of the solution is given. In Section 4, the propagation of chaos for this model is also analyzed. In Section 5, the EM scheme is established and the strong convergence rate is given.
@@ 2 Preliminaries
[…]
$\mathcal{P}_{q}(\mathbb{R}^{d})=\bigg{\{}\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)<\infty\bigg{\}},$
which is a Polish space under the $L^{q}$-Wasserstein distance [30, Definition 6.1]. Then, for $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, define
$\mathcal{W}_{q}(\mu)=\bigg{(}\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)\bigg{)}^{1/q}$
For $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, we define the Wasserstein distance on $\mathcal{P}_{q}(\mathbb{R}^{d})$ by
$\mathbb{W}_{q}(\mu,\nu)=\bigg{(}\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{q}\pi({\rm d}x,{\rm d}y)\bigg{)}^{1/q},$
where $\Pi(\mu,\nu)$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$. In other words, $\pi\in\Pi(\mu,\nu)$ is a probability measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that $\pi(\cdot\times\mathbb{R}^{d})=\mu(\cdot)$ and $\pi(\cdot\times\mathbb{R}^{d})=\nu(\cdot)$. For $11/2<H<$ 1, we put
@@
and recall that for $s,t>0$,
$\int_{0}^{t}\int_{0}^{s}\phi(u,v){\rm d}u{\rm d}v=R_{H}(t,s).$
Let $\mathcal{S}(\mathbb{R})$ be the Schwartz space of rapidly decreasing smooth functions on $R$, and if $b\in\mathcal{S}(\mathbb{R})$, denote
$\parallel f\parallel_{H}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)f(t)\phi(s,t){\rm d}s{\rm d}t<\infty.$
If we equip $\mathcal{S}(\mathbb{R})$ with the inner product
$\langle f,g\rangle_{H}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)g(t)\phi(s,t){\rm d}s{\rm d}t,\ \ f,g\in\mathcal{S}(\mathbb{R}),$
then the completion of $\mathcal{S}(\mathbb{R})$, denoted by $L_{\phi}^{2}(\mathbb{R})$,
@@
[32, Definition 3.5.1] Let $g\in L_{\phi}^{2}(\mathbb{R})$. The $\phi$-derivative of a stochastic variable $F\in L^{p}$ in the direction of $\Phi_{g}$ is defined by
$D_{\Phi_{g}}F(\omega)=\lim_{\delta\to 0}\frac{1}{\delta}\bigg{\{}F\bigg{(}\omega+\delta\int_{0}^{\cdot}(\Phi_{g})(u){\rm d}u\bigg{)}-F(\omega)\bigg{\}},$
if the limit exists in $L^{p}$. Furthermore, if there is a process $(D_{s}^{\phi}F,s\geq 0)$ such that
$D_{\Phi_{g}}F=\int_{0}^{\infty}D_{s}^{\phi}Fg_{s}{\rm d}s\ \ \ {\rm almost\ surely},$
for all $g\in L_{\phi}^{2}(\mathbb{R})$, then $F$ is said to be $\phi$-differentiable.
For SDEs driven by fractional Brownian motion
(2.2) ${\rm d}X_{t}=G(t,X_{t}){\rm d}t+F(t,X_{t}){\rm d}B_{t}^{H}.$
Let $C^{2,1}(\mathbb{R}\times\mathbb{R}_{+};\mathbb{R})$ be the family of all real-valued functions $V(x,t)$ defined on $\mathbb{R}\times\mathbb{R}_{+}$, which is twice continuously differentiable in $x$ and once in $t$. Referring to the Itô formula for fractional Brownian motion given by [32, 33],
@@
or formally,
${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G_{t,X_{t}}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
@@
In addition, [34], [35] also discussed the fractional Itô formula. Now, we are going to give the fractional Itô formula as follows.
Remark 2.1.
(Fractional Itô formula) Assume that the stochastic process $X_{t}$ follows (2.2) with $11/2<H<$ 1. We have
@@
$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})$
$+H(2H-1)\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}\int_{0}^{s}(s-r)^{2H-2}F^{2}(r,X_{r}){\rm d}r\bigg{]}{\rm d}s+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}$
$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})+H\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}s^{2H-1}F^{2}(s,X_{s})\bigg{]}{\rm d}s$
$+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}.$
Since the kernel function $\phi$ is defined only for $11/2<H<$ 1, otherwise, $\phi$ has no meaning. We assume that $11/2<H<$ 1 throughout this paper. In this paper, we consider the following $d$-dimensional McKean-Vlasov NSDDEs driven by fractional Brownian motion as follows:
(2.5) $\left\{\begin{array}[]{ll}{\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}_{X_{t}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}}){\rm d}B_{t}^{H},\ \ t>0,\\ X_{t}=\xi=\{\xi(t):-\tau\leq t\leq 0\}\in L_{\mathcal{F}_{0}}^{p}([-\tau,0];\mathbb{R}^{d}),\ p\geq\theta\geq 1\end{array}\right.$
where $b:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$, $\mathcal{L}_{X_{t}}$ denotes the law of $X_{t}$. Throughout the paper, we shall denote $C$ by a generic positive constant, whose value may change from line to line. For any $x,y,\overline{x},\overline{y}$, we impose the following assumptions.
@@
$D(0)=0$, and there exist positive constants $\lambda<1$ such that
$|D(x)-D(\overline{x})|\leq\lambda|x-\overline{x}|.$
@@
There exist positive constants $l\geq 1$ and $K_{2},K_{3}$ such that
$\big{\langle}x-D(y)-\overline{x}+D(\overline{y}),b(x,y,\mu)-b(\overline{x},\overline{y},\nu)\big{\rangle}\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
@@
$|b(x,y,\mu)-b(\overline{x},\overline{y},\nu)|\leq K_{3}[(1+|x|^{l}+|\overline{x}|^{l}+|y|^{l}+|\overline{y}|^{l})(|x-\overline{x}|+|y-\overline{y}|)+\mathbb{W}_{q}(\mu,\nu)],$
for $x,y,\overline{x},\overline{y}\in\mathbb{R}^{d}$, $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$.
@@
There exist positive constants $K_{4},K_{5}$ such that
$\parallel\sigma(\mu)-\sigma(\nu)\parallel\leq K_{4}\mathbb{W}_{q}(\mu,\nu),$
and,
$|b(0,0,\mu)|\vee\parallel\sigma(\mu)\parallel\leq K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0})),$
for $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$. Then, for initial experience distribution $\delta_{0}$
$|b(0,0,\delta_{0})|\vee\parallel\sigma(\delta_{0})\parallel\leq K_{6}.$

### v2

# Strong convergence of tamed theta scheme for superlinearly growing McKean-Vlasov NSDDEs driven by fractional Brownian motionsthanks: Supported by Jiangxi Provincial Natural Science Foundation(No., 20252BAC250002, GJJ2404201)

@@ Abstract
In this article, we study the McKean-Vlasov neutral stochastic differential delay equations driven by fractional Brownian motion with super-linearly growing coefficients, where the Hurst exponent $H\in(1/2,1)$. The existence and uniqueness of the exact solution were shown by the Picard iteration. Besides, we propose a tamed theta Euler-Maruyama scheme for this equation, analyzed the moment boundness and propagation of chaos etc. Moreover, the convergence rate of the numerical scheme is established.
Keywords : McKean-Vlasov NSDDEs; super-linearly growing coefficients; fractional Brownian motion; strong convergence rate; tamed theta EM scheme
@@
Stochastic differential equations (SDEs) whose coefficients depend on the law of the solution (also called McKean-Vlasov SDEs), were first investigated by McKean [1], whose work drew inspiration from Kac’s programme in Kinetic Theory [2]. Since then, McKean-Vlasov SDEs have evolved into a major area of research in stochastic analysis. To approximate the McKean-Vlasov SDE limit, Bossy and Talay [3, 4] proposed a system of $n$ weakly interacting particles—replicating the statistical behavior of the limit system, and systematically studied the numerical approximation of its solutions. Building on this foundation, the existence and uniqueness of McKean-Vlasov SDEs were discussed in [5, 6], while the well-posedness and the propagation of chaos for McKean-Vlasov SDEs were investigated by Lacker [7]. Given the rarity of explicit solutions for McKean-Vlasov SDEs, numerical approximation methods have become indispensable. A variety of schemes have been developed to handle different coefficients behaviors, for example the implicit Euler-Maruyama (EM) scheme [8, 9], the adaptive EM scheme [10, 11], the tamed EM scheme [12, 13], and truncated EM scheme [14].
In recent years, research attention has increasingly shifted to McKean-Vlasov SDEs driven by fractional Brownian motion—a generalization of standard Brownian motion with unique self-similarity and long-memory properties, making it well-suited for modeling phenomena in atmospheric sciences, financial engineering, and other fields. These equations typically take the form:
@@
where the coefficients $b:\mathbb{R}^{d}\times\mathcal{P}_{\theta}(\mathbb{R}^{d})\to\mathbb{R}^{d},\sigma:\mathcal{P}_{\theta}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ are Borel-measurable, here, $\mathcal{P}_{\theta}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite $\theta$-th moment. $\mathcal{L}(X_{t})$ is the law of $X$ at $t$ and $B^{H}=\{B_{t}^{H},t\in[0,T]\}$ is a fractional Brownian motion with Hurst exponent $H\in(0,1)$ defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. As a centered Gaussian process, $B_{H}$ has the covariance function
(1.2) $R_{H}(t,s)=\mathbb{E}(B_{t}^{H}B_{s}^{H})=\frac{1}{2}(t^{2H}+s^{2H}-|t-s|^{2H}),\ \ \forall t,s\in[0,T].$
Notably, when $H=1/2$, the fractional Brownian motion $\{B_{t}^{H}\}_{t\geq 0}$ reduces to a standard Brownian motion. The well-posedness of such equations was first proven by Fan et al.[15]. Subsequent research have further expanded the theoretical framework of this model. For example, under the assumption that the coefficients satisfy globally Lipschitz conditions, He et al. [16] established the propagation of chaos for the model described in (1.1) and analyzed the strong convergence of the EM method.
@@
A critical gap in current literature is the lack of studies on McKean-Vlasov SDEs driven by fractional Brownian motion with super-linearly growing coefficients. One notable exception is Shen et al.[17] who showed that under specific averaging conditions, solutions to such equations can be approximated by solutions of an associated averaged equation. More recently, Wang et al. [18] and Gao et al. [19] focused on the EM scheme for delayed McKean-Vlasov SDEs driven by fractional Brownian motion, but their work only considered super-linear growth in the delay term of the drift coefficient, not the coefficient itself.
To capture the influence of past states on current dynamics, stochastic delay differential equations (SDDEs) and neutral stochastic differential delay equations (NSDDEs) are essential tools in stochastic calculus, with broad applications in modeling natural and engineered systems, see for example [20, 21, 22]. A key challenge arises when NSDDEs exhibit super-linear growth: the classical EM scheme often fails to converge. To address this, a tamed EM scheme was introduced to handle super-linearly growing coefficients. For instance, Deng et al. [23] proposed two types of explicit tamed EM schemes for NSDDEs with super-linearly growing drift and diffusion coefficients and proved strong convergence for each. Cui et al. [24] studied the tamed EM methods to approximate the solutions of McKean-Vlasov NSDDEs with super-linear drift, establishing the convergence rate. Additional insights into tamed EM schemes for NSDDEs can be found in [22, 25]. Notably, Tan and Yuan [26] proposed a modified tamed theta EM scheme and demonstrated the convergence of the numerical solutions, providing a key motivation for the present work.
Drawing inspiration from the existing literatures, we extend the model (1.1) to accommodate coefficients with super-linear growth. We propose a tamed theta EM scheme for the McKean-Vlasov NSDDEs, which takes the following form:
@@
where $\tau>0$ denotes the delay.
This paper’s main contributions are outlined below:
  • Unlike the prior work (e.g., [18, 19]) that only considers super-linear growth in the delay term, we analyze McKean-Vlasov equations driven by fractional Brownian motion where the drift coefficient itself exhibits super-linear growth. We prove the core properties of such equations, representing a significant advance in this field.
  • By a delicate analysis involving the fractional Itô formula and intricate inequality techniques, we investigate the properties of the solution, the propagation of chaos in associated particle systems, and the convergence of the proposed numerical scheme.
The rest of the paper is organized as follows. Section 2 presents the mathematical preliminaries, including key definitions and lemmas for McKean-Vlasov NSDDEs with super-linear growth driven by fractional Brownian motion. Section 3 establishes the existence and uniqueness of the solution. In Section 4, we analyze the propagation of chaos in the particle system associated with our model. Finally, Section 5 introduces the tamed theta EM scheme and derives its strong convergence rate.
@@ 2 Preliminaries
Throughout this paper, let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space with a filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ satisfying the usual conditions (i.e., it is increasing and right continuous while $\mathcal{F}_{0}$ contains all $\mathbb{P}$-null sets). We use $\lvert\cdot\rvert$ and $\langle\cdot,\cdot\rangle$ for the Euclidean norm and inner product, respectively, and let $a\wedge b:=\min(a,b)$ and $a\vee b:=\max(a,b)$. Let $\tau>0$ be a constant and denote by $\mathcal{C}([-\tau,0];\mathbb{R}^{d})$ the space of all continuous functions $\varphi$ from $\tau,0]$ to $\mathbb{R}^{d}$ with the norm $\parallel\varphi\parallel=\sup_{-\tau\leq t\leq 0}|\varphi(t)|$. For $p>0$, $L_{\mathcal{F}_{0}}^{p}(\Omega;\mathbb{R}^{d})$ represents the family of $\mathcal{F}_{0}$-measurable random variable with $\mathbb{E}|\cdot|^{p}<\infty$. Let $\mathcal{P}_{\theta}(\mathbb{R}^{d}),\theta\geq 1$ be the space of probability measures on $\mathbb{R}^{d}$ with finite $\theta$-th moment. Formally, it is defined as
$\mathcal{P}_{\theta}(\mathbb{R}^{d})=\bigg\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{\theta}\mu({\rm d}x)<\infty\bigg\},$
where $\mathcal{P}(\mathbb{R}^{d})$ denotes the family of all probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, here, $\mathcal{B}(\mathbb{R}^{d})$ is the Borel $\sigma$-field over $\mathbb{R}^{d}$. For any $\mu\in\mathcal{P}_{\theta}(\mathbb{R}^{d})$, define the $\theta$-th moment norm of $\mu$ as
$\mathcal{W}_{\theta}(\mu)=\bigg(\int_{\mathbb{R}^{d}}|x|^{\theta}\mu({\rm d}x)\bigg)^{1/\theta}.$
Next, for two probability measures $\mu,\nu\in\mathcal{P}_{\theta}(\mathbb{R}^{d})$, the Wasserstein distance between $\mu$ and $\nu$ is defined as
$\mathbb{W}_{\theta}(\mu,\nu)=\bigg(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{\theta}\pi({\rm d}x,{\rm d}y)\bigg)^{1/\theta},$
where $\Pi(\mu,\nu)$—the set of coupling measures between $\mu$ and $\nu$, consists of all probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ whose marginal measures are $\mu$ and $\nu$, respectively. Precisely, a measure $\pi\in\Pi(\mu,\nu)$ satisfies $\pi(A\times\mathbb{R}^{d})=\mu(A)$ and $\pi(\mathbb{R}^{d}\times A)=\nu(A)$ for every Borel set $A\subseteq\mathbb{R}^{d}$. The space $\mathcal{P}_{2}(\mathbb{R}^{d})$ is a Polish space when equipped with the $L^{2}$-Wasserstein distance [27, Definition 6.1].
Let $\{B_{t}^{H}\}_{t\in[0,T]}$ be a $d$-dimensional fractional Brownian motion with Hurst exponent $H\in(0,1)$. For Hurst exponents $1/2<H<1$, we define the kernel function
@@
and recall that for $s,t>0$, this kernel satisfies the integral identity
$\int_{0}^{t}\int_{0}^{s}\phi(u,v){\rm d}u{\rm d}v=R_{H}(t,s),$
where $R_{H}(t,s)$ is defined as in (1.2). Let $\mathcal{S}(\mathbb{R})$ be the Schwartz space (space of rapidly decreasing smooth functions) on $R$. For any $f\in\mathcal{S}(\mathbb{R})$, we define its squared norm with respect to the kernel $\phi$ as
$\|f\|_{H}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)f(t)\phi(s,t){\rm d}s{\rm d}t<\infty,$
where $\phi$ is the kernel defined in (2.1). Next, we equip $\mathcal{S}(\mathbb{R})$ with an inner product induced by $\phi$ as
$\langle f,g\rangle_{H}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)g(t)\phi(s,t){\rm d}s{\rm d}t,\ \ f,g\in\mathcal{S}(\mathbb{R}).$
The completion of $\mathcal{S}(\mathbb{R})$ with respect to this inner product is denoted by $L_{\phi}^{2}(\mathbb{R})$. For the half line $\mathbb{R}_{+}:=[0,+\infty)$, we further define the space $L_{\phi}^{2}(\mathbb{R}_{+})$ as the set of measurable functions $g:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that their squared $\phi$-norm is finite:
@@
Here, $L_{\phi}^{2}(\mathbb{R}_{+})$ can be understood as the natural restriction of $L_{\phi}^{2}(\mathbb{R})$ to functions supported on $\mathbb{R}_{+}$, and it is also a Hilbert space under the inner product
$\langle g_{1},g_{2}\rangle_{\phi}:=\int_{0}^{\infty}\int_{0}^{\infty}g_{1}(s)g_{2}(t)\phi(s,t){\rm d}s{\rm d}t.$
@@
[28, Definition 3.5.1] Let $g\in L_{\phi}^{2}(\mathbb{R})$, and $\Phi_{g}$ denote a random variable induced by $g$, i.e., $\Phi_{g}=\int_{\mathbb{R}}g(s){\rm d}B_{s}^{H}$. For $F\in L^{p}$, the $\phi$-derivative of $F$ in the direction of $\Phi_{g}$ is defined as
$D_{\Phi_{g}}F(\omega)=\lim_{\delta\to 0}\frac{1}{\delta}\bigg\{F\bigg(\omega+\delta\int_{0}^{\cdot}g(u){\rm d}u\bigg)-F(\omega)\bigg\},$
provided the limit exists in $L^{p}$. Furthermore, if there is a process $(D_{s}^{\phi}F,s\geq 0)$ such that for all $g\in L_{\phi}^{2}(\mathbb{R})$
$D_{\Phi_{g}}F=\int_{0}^{\infty}D_{s}^{\phi}F\cdot g_{s}{\rm d}s\ \ \ {\rm almost\ surely},$
then $F$ is said to be $\phi$-differentiable.
We denote $L_{\phi}(0,T)$ as the set of stochastic processes $F$ on $[0,T]$ that satisfy the following properties: $F\in L_{\phi}(0,T)$ if and only if $\mathbb{E}\|F\|_{H}^{2}<\infty$, $F$ is $\phi$-differentiable, the trace $D_{s}^{\phi}F_{t}$ exists for all $0\leq s,t\leq T$, $\mathbb{E}\int_{0}^{T}\int_{0}^{T}|D_{s}^{\phi}F_{t}|^{2}\,ds\,dt<\infty$, and for each sequence of partitions $(\pi_{n},n\in\mathbb{N})$ such that $\|\pi_{n}\|\to 0$ as $n\to\infty$, it holds that
$\sum_{i,j=0}^{n-1}\mathbb{E}\left[\int_{t_{i}^{(n)}}^{t_{i+1}^{(n)}}\int_{t_{j}^{(n)}}^{t_{j+1}^{(n)}}|D_{s}^{\phi}F_{t_{i}^{(n)}}^{\pi}D_{t}^{\phi}F_{t_{j}^{(n)}}^{\pi}-D_{s}^{\phi}F_{t}D_{t}^{\phi}F_{s}|dsdt\right]\to 0,$
and
$\mathbb{E}\|F^{\pi}-F\|_{H}^{2}\to 0,n\to\infty,$
where $\pi_{n}:0=t_{0}^{(n)}<t_{1}^{(n)}<\cdots<t_{n-1}^{(n)}<t_{n}^{(n)}=T.$
Consider the $d$-dimensional SDEs driven by a $d$-dimensional fractional Brownian motion $B_{t}^{H}$ with Hurst exponent $H\in(1/2,1)$ as follows
(2.2) ${\rm d}X_{t}=G(t,X_{t}){\rm d}t+F(t,X_{t}){\rm d}B_{t}^{H},$
where $G:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $F:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$. Let $C^{2,1}([0,T]\times\mathbb{R}^{d};\mathbb{R})$ denote the space of real-valued functions $V(t,x)$ that are twice continuously differentiable in $x\in\mathbb{R}^{d}$ and once continuously differentiable in $t\in[0,T]$. Based on the fractional Itô formula derived in [28, 29], the differential form of $V(t,X_{t})$ takes the form:
@@
where $\frac{\partial V}{\partial X}$ denotes the gradient, $\frac{\partial^{2}V}{\partial X^{2}}$ is the Hessian matrix. Formally, this can also be expressed by the kernel $\phi$ as
@@
Additional discussions on the fractional Itô formula can be found in [30, 31]. We summarize the result below.
Lemma 2.1.
(Fractional Itô formula) Let $X_{t}$ be the solution of (2.2) with $1/2<H<1$, and assume that $\mathbb{E}\bigg[\int_{0}^{T}\left\lvert F(s,X_{s})D_{s}^{\phi}\int_{0}^{s}F(u,X_{u}){\rm d}B_{u}^{H}\right\rvert{\rm d}s\bigg]<\infty$, $\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s})\in L_{\phi}(0,T)$ for $V\in C^{2,1}([0,T]\times\mathbb{R}^{d};\mathbb{R})$. Then we have
@@
$=$ $V(0,X_{0})+\int_{0}^{t}\bigg[\frac{\partial V(s,X_{s})}{\partial s}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})$
$+H(2H-1)\text{{\rm tr}}\left(\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}\int_{0}^{s}(s-r)^{2H-2}F(r,X_{r})F^{T}(r,X_{r}){\rm d}r\right)\bigg]{\rm d}s$
$+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H},$
where $F^{T}$ is the transpose of $F$, and $\text{{\rm tr}}(\cdot)$ denotes the matrix trace.
Since the kernel function $\phi$ is only valid for $1/2<H<1$, we restrict our analysis to Hurst exponents satisfying $1/2<H<1$ throughout this paper. In this work, we focus on the following $d$-dimensional McKean-Vlasov NSDDEs driven by fractional Brownian motion as follows:
(2.5) $\left\{\begin{array}[]{ll}{\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}_{X_{t}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}}){\rm d}B_{t}^{H},\ \ t>0,\\ X_{t}=\xi=\{\xi(t):-\tau\leq t\leq 0\}\in L_{\mathcal{F}_{0}}^{{p}}([-\tau,0];\mathbb{R}^{d}),\ {p}\geq 1\end{array}\right$.
where $b:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}^{d\times d}$, $D:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $\mathcal{L}_{X_{t}}$ denotes the law of $X_{t}$. Throughout the paper, we shall denote by $C$ a generic positive constant that may depend on $T,H,\xi,l,\lambda,K_{i}$ etc., but independent of $n,m,s,t,N,\Delta$. The value may change from line to line.
@@
$D(0)=0$, and there exists a constant $0<\lambda<1$ such that
$|D(x)-D(\overline{x})|\leq\lambda|x-\overline{x}|,$
for $x,\overline{x}\in\mathbb{R}^{d}$.
@@
There exist positive constants $l\geq 1$, $K_{2}$ and $K_{3}$ such that
$\big\langle x-D(y)-\overline{x}+D(\overline{y}),b(x,y,\mu)-b(\overline{x},\overline{y},\nu)\big\rangle\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+\mathbb{W}_{2}^{2}(\mu,\nu)),$
@@
$|b(x,y,\mu)-b(\overline{x},\overline{y},\nu)|\leq K_{3}[(1+|x|^{l}+|\overline{x}|^{l}+|y|^{l}+|\overline{y}|^{l})(|x-\overline{x}|+|y-\overline{y}|)+\mathbb{W}_{2}(\mu,\nu)],$
for $x,y,\overline{x},\overline{y}\in\mathbb{R}^{d}$, $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
@@
There exist a positive constant $K_{4}$ such that
$\parallel\sigma(\mu)-\sigma(\nu)\parallel\leq K_{4}\mathbb{W}_{2}(\mu,\nu).$
And for the initial experience distribution $\delta_{0}$, there is a positive constant $K_{5}$ such that
$|b(0,0,\delta_{0})|\vee\parallel\sigma(\delta_{0})\parallel\leq K_{5}.$